OpenStudy (anonymous):

Domain of y = (x+3) / 4- sqrt(x^2-9)

OpenStudy (anonymous):

\[y=x+3\div4-\sqrt{x ^{2}-9}\]

OpenStudy (anonymous):

basically everything x can be >= 3

OpenStudy (anonymous):

else itwould be imaginary... unless you want that

OpenStudy (anonymous):

Isn't there a restriction at +- 5?

OpenStudy (anonymous):

Why xcan't be x<=-3?

OpenStudy (anonymous):

That's what I'm saying... can't it go to -infintiy?

OpenStudy (anonymous):

(-infinity, -3)U(-3,-5), (3,5)U(5,infinity) ?

OpenStudy (anonymous):

\[\left( -\infty,-5 \right)\cup(-5,-3]\cup[3,5)\cup(5,\infty)\]